OpenStudy (anonymous):

If the midpoint of a segment is (-2,0) and one endpoint is (3,5) what is the other endpoint?

OpenStudy (anonymous):

-7 and -5

OpenStudy (anonymous):

so to go from (-2,0) to (3,5), you need to move 5 units to the right and 5 units up. We need to do the same thing to get the other endpoint- moving up 5, right 5 from (3,5) gives us (8,10).

OpenStudy (anonymous):

GT, Can you explain how you got that?

OpenStudy (mertsj):

You went 5 to the left of (3,5) to get to the midpoint so go 5 more to the left to get to the other end. Five to the left of -2 is -7. Also you went down 5 to get from the endpoint to the midpoint so go down 5 more to get to the other end. Five down from 0 is -5 so the other endpoint must be (-7,-5)

OpenStudy (anonymous):

MertsJ - your answer is not right. Midpoint of two points (x1, y1) and (x2, y2) is (x1+x2)/2, (y1+y2)/2. So, here you are given one endpoint and the midpoint. (x1+3)/2 = -2 (y1+5)/2 = 0 Solve for x1 and y1.

OpenStudy (anonymous):

Sorry. Your answer is right MertsJ!

OpenStudy (mertsj):

It's the same as yours. GT 0 Good Answer -7 and -5

OpenStudy (anonymous):

That clears up a lot for me. I have another question like this. I going to try it 1st and I'll post my answer. Can you tell me if it's right?

OpenStudy (anonymous):

Yes. Medal for me? :)

OpenStudy (anonymous):

You got it.

OpenStudy (anonymous):

Q:If the midpoint of a segment is (-2, 4) and one endpoint is (1, 2), find the other endpoint. A: I got (-5,6)

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

Awesome. You've been a great help.

OpenStudy (anonymous):

You have been an awesome learner!

OpenStudy (anonymous):

Thanks. Appreciate that.